shift summary, day 02/03/00


Subject: shift summary, day 02/03/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:51:18 CST


Shift Summary for day shift, 02/03/00

On Shift: Pat T., Ed B.

Hours of beam delivered at all intensities: 7.5
Hours of beam delivered at nominal intensity: 7.5
Hours of KAMI data: 7.5

Accounting of lost time: Minor accelerator problems, water chiller failure

===============================================================

Runs ended this shift:
15751 KAMI_MINBIAS BAG west beam hole, 20" Be
15752 KAMI_MINBIAS BAG east beam hole, 20" Be
15753 KAMI_MINBIAS 0" Be
15755 KAMI_MINBIAS 0" Be
15756 KAMI_MINBIAS 0" Be
15757 KAMI_MINBIAS 0" Be

===============================================================

Tapes brought to FCC this shift:

Accesses during shift: 1) move BAG to east beam hole
                       2) reset ET after chiller shuts off

===============================================================

Description of Shift: KAMI minbias running at targeting angle
                      of 15 mrad

run spill time comment
----------------------------------------------------
15751 09:05 Stop run 15751.
             09:20 Bob H. and Anatoli go on access to move BAG
                       from west beam hole to east beam hole.

15752 09:30 Start run 15752.
             10:00 Bob H. notices all swic beam data has disappeared
                       along with counts in the accidental counter.
                       Beam is off target due to an LCW trip.
             10:25 Beam back on target.
             11:00 Stop run 15752.

15753 11:15 Start run 15753.
             11:20 Increasing temperature and ET alarms indicate
                       that the water chiller has tripped off again.
                       Duty mechanic shows up promptly to reset it.
                       S.T. and J.L. go on acces to reset ET. All other
                       systems are fine.
             11:30 Stop run 15753.

15755 12:00 Start run 15755.
             14:00 Stop run 15755.

15756 14:05 Start run 15756.
             14:35 Notice that no C07 triggers are getting through
                       L3. Run is stopped and namelist is modified.

15757 14:40 Start run 15757.
             14:55 Beam gone, MI extraction problems.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:51:23 CST